# Smallest Suffixient Sets: Effectiveness, Resilience, and Calculation

### v4

@@ 4.1 Appending and prepending symbols
Cases 1 and 2 induce at most one new super-maximal right-extension in total for all possible $xc$, namely the longest right-extension in $wc$ that is a suffix of $wc$. For Case 3, consider a fixed $a\in\Sigma$. For all the increasing-length suffixes $x_{1}c,x_{2}c,\dots,x_{t}c$ of $wc$ that became right-extensions together with $x_{1}a,x_{2}a,\dots,x_{t}a$, one can see that the latter form a chain of suffixes of $x_{t}a$. Hence, we only have one possible new super-maximal right-extension ending with $a$, namely $x_{t}a$. Observe that the chain of suffixes $x_{1}a,x_{2}a,\dots,x_{t}a$ is unique: if the suffix $x$ is always followed by $a$, any suffix $y$ of $x$ is either right-maximal in $w$ (and $y$ falls within Case 2), or it is always followed by an $a$ (because $x$ is always followed by an $a$), i.e. $y=x_{i}$ for some $i\in[1\mathinner{.\,.}t]$. ∎
Lemma 3
@@
Proof
The lower bound follows from Lemma 1. For the upper bound, let cxacxa be the smallest prefix of $cw$ that is not a right-extension of $w$, but is a right-extension of $cw$ (if it exists). This implies that there exists $b\neq a$ such that $cxb\in\mathcal{F}_{w}$ and $cxa\not\in\mathcal{F}_{w}$ (otherwise, cxacxa would be a right-extension of $w)$, so no prefix of $cw$ of length $|cxa|$ or more is right-maximal, and thus no prefix longer than cxacxa can be a right-extension. By the minimality of cxacxa, all prefixes of $cw$ shorter than cxacxa either are not right-extensions of $cw$, or are right-extensions of both $w$ and $cw$. Therefore, cxacxa together with some cxbcxb appearing in $w$, are the only possible new right-extensions in $cw$ with respect to ww.∎
Lemmas 2 and 3 yield a remarkably simple algorithms to compute smallest suffixient sets, which we detail in Section 7.
By letting $c=\text{\tt\textdollar}\not\in\mathcal{F}_{w}$ in Lemma 2, we relate $\chi$ to $\mathtt{sre}$ (note that $\chi$ is always at least $\mathtt{sre}+1$ because of the new super-maximal extension ending with \$). This makes clear the relation between Combinatorics on words [21] with suffixient sets, via the common notion of right-special factors (which we call here right-maximal substrings).
Corollary 1
@@
4.2 Arbitrary edit operations and rotations
Now we study how much $\mathtt{sre}(w)$ can vary upon edit operations at arbitrary positions and rotations. We will use the following famous string family.
Definition 6
Lemma 5
Proof
Let $w[1\mathinner{.\,.}n]$ be a binary de Bruijn string of order $k$. By definition, $w$ contains every binary string of length $k$ as a substring exactly once. As all the possible pairs of strings $x\texttt{a}$ and $x\texttt{b}$ of length $k$ appear in $w$, it follows that all the strings in $\mathcal{F}_{w}(k)$ are right-extensions. Moreover, each $x\texttt{a}$ and $x\texttt{b}$ of length $k$ are super-maximal right-extensions: otherwise, there would exist some $c\in\{\texttt{a},\texttt{b}\}$ such that $cx\texttt{a}$ and $cx\texttt{b}$ are both substrings of $w$, which raises a contradiction since the $k$-length string $cx$ cannot appear twice in $w$. Moreover, there are no right-maximal strings of length $k$ or greater; hence, there are no right-extensions of length greater than $k$. It follows that $\mathtt{sre}(w)=|\mathcal{F}_{w}(k)|=2^{k}=\Omega(n)$. ∎
The following lemma uses the de Bruijn family to show that $\mathtt{sre}$ can grow by $\Omega(\log n)$ upon arbitrary edit operations and rotations.
Lemma 6
Let $w=\texttt{a}^{k}\texttt{b}\texttt{a}^{k-2}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1}\in\mathtt{dB}(k)$ be the lexicographically smallest binary de Bruijn sequence of order $k$ [11, 12]. It holds:
@@ 4.2 Arbitrary edit operations and rotations
    (Ins) $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-2$ if $w^{\prime}=\texttt{a}^{2k-2}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1}$,
@@
    (Sub) $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-3$ if $w^{\prime}=\texttt{a}^{k}\texttt{b}\texttt{a}^{k-2}\texttt{b}x\texttt{a}\texttt{b}^{k-1}\texttt{c}\texttt{a}^{k-1}$,
@@
    (Del) $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-4$ if $w^{\prime}=\texttt{a}^{k}\texttt{b}\texttt{a}^{k-2}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{c}\texttt{a}^{k-1}$,
@@
    (Rot) $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-2$ if $w^{\prime}=\texttt{b}\texttt{a}^{k-2}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{2k-1}$.
Proof
Observe that in each claim, $w$ is obtained after performing a string operation on the corresponding $w^{\prime}$: in Claim 1, $w\in{\tt ins}_{\Sigma}(w^{\prime})$; in Claim 2, $w\in{\tt sub}_{\Sigma}(w^{\prime})$; in Claim 3, $w\in{\tt del}_{\Sigma}(w^{\prime})$; in Claim 4, $w\in\mathcal{R}(w^{\prime})$. We prove each claim separately by comparing the super-maximal extensions of $w^{\prime}$ before and after performing the string operation on $w^{\prime}$ that yields $w$, for which $\mathtt{sre}(w)=2^{k}$ by Lemma 5.
For Claim 1, note that $\mathtt{sre}(w^{\prime})$ is the same as $\mathtt{sre}(\texttt{a}^{k}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1})$, as prepending the character a multiple times to this string to obtain $w^{\prime}$ never increases $\mathtt{sre}$; it only updates the super-maximal extension $\texttt{a}^{k}$ to $\texttt{a}^{k+1}$ and $\texttt{a}^{k-1}\texttt{b}$ to $\texttt{a}^{k}\texttt{b}$, and so on. For simplicity, we let $w^{\prime}=\texttt{a}^{k}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1}$. The string $w^{\prime}$ does not contain substrings of length $k$ of the form $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}$ for $i\in[1\mathinner{.\,.}k-2]$, nor the substring $\texttt{b}\texttt{a}^{k-2}\texttt{b}$. Note that for each of these substrings $y\in\mathcal{F}_{w}(k)$ with $y\not\in\mathcal{F}_{w^{\prime}}(k)$, the other corresponding right-extension $y^{\prime}$ in $w$ sharing a length k1k-1 prefix with $y$ is not a right-extension in $w^{\prime}$. Moreover, note that all the suffixes of length k1k-1 of these $y$ are not suffixes of one another, nor of the length k1k-1 suffixes of any of the substrings $y^{\prime}$ in $w^{\prime}$. Hence, all k1k-1 length binary strings still appear in $w^{\prime}$ as the suffix of some length $k$ substring that remains a right-extension in $w^{\prime}$, and hence, super-maximal extensions of $w^{\prime}$ have to be of length at least $k$. As each string of length $k$ appearing in $w^{\prime}$ is unique, there are no super-maximal extensions of length greater than $k$. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)$ because we are losing k1k-1 pairs of super-maximal extensions of length $k$ with respect to $w$. It follows that by inserting the b in $w^{\prime}$ to yield $w$, $\mathtt{sre}$ increases by 2k22k-2.
For Claim 2, note that exactly $k$ substrings of length $k$ are lost when substituting the last b of $w$ by c: those of the form $\texttt{b}^{i}\texttt{a}^{k-i}$ with $i>0$. This means that substrings ending in $\texttt{b}^{i}\texttt{a}^{k-i-1}$ with $0<i<k$ are not right-maximal in $w^{\prime}$, hence, 2(k1)2(k-1) super-maximal extensions are lost. Moreover, $\texttt{b}^{k-2}$ is still a right-maximal substring, since $\texttt{b}^{k-1}$ and $\texttt{b}^{k-2}\texttt{c}$ occur in $w^{\prime}$. Observe that only $\texttt{b}^{k-2}\texttt{c}$ is a super-maximal extension, while $\texttt{b}^{k-1}$ is a suffix of $\texttt{a}\texttt{b}^{k-1}$. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)+1$ and $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-3$.
For Claim 3, the analysis is similar to Claim 2, but in $w^{\prime}$, $\texttt{b}^{k-1}$ remains as a super-maximal extension. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)+2$ and $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-4$.
For Claim 4, the analysis is similar to Claim 1, but in $w^{\prime}$, $\texttt{b}\texttt{a}^{k-2}\texttt{b}$ appears, while $\texttt{a}^{k-1}\texttt{b}$ does not. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)$ and $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-2$. ∎
4.3 Reversals
We now show that $\mathtt{sre}$ can grow by $\Omega(\sqrt{n})$ upon string reversals.
Lemma 7
Let $k>0$. Let $w_{k}=\prod_{i=0}^{k-1}\texttt{c}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{i}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\textdollar}_{i}$ on the alphabet $\Sigma=\{\texttt{a},\texttt{b},\texttt{c}\}\cup\bigcup_{i\in[0\mathinner{.\,.}k-1]}\{\text{\tt\#}_{i},\text{\tt\textdollar}_{i}\}$. It holds $\mathtt{sre}(w_{k})-\mathtt{sre}(w_{k}^{R})=k-1$.
Proof
Observe that by construction, any substring of $w_{k}$ containing $\text{\tt\#}_{i}$ or $\text{\tt\textdollar}_{i}$ is not right-maximal, as these symbols are unique. Hence, the right-extensions of $w_{k}$ cannot cross from one side to the other side of those special delimiters. Moreover, substrings of the form $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}$ for $i\in[0\mathinner{.\,.}k-1]$ appear exactly twice in $w_{k}$ and their right-extensions are super-maximal. By looking at the structure of the string $w_{k}$ and carefully analyzing its right-extensions, one can verify that the super-maximal right-extensions of $w_{k}$ are the following:
@@ 4.3 Reversals
    $\texttt{b}\texttt{a}^{k-1}$ and c
@@
    $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{i}$ and $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\textdollar}_{i}$ for $i\in[0\mathinner{.\,.}k-1]$,
@@
    $\texttt{c}\texttt{a}^{i}$ and $\texttt{c}\texttt{a}^{i-1}\texttt{b}$ for $i\in[1\mathinner{.\,.}k-1]$,
@@
    $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}$ for $i\in[1\mathinner{.\,.}k-1]$.
This sums to a total of 5k15k-1 super-maximal extensions in $w_{k}$. In the reversed string $w_{k}^{R}=\prod_{i=0}^{k-1}\text{\tt\textdollar}_{k-i-1}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{k-i-1}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\texttt{c}$, we have instead:
@@
    $\texttt{b}\texttt{a}^{k-1}$ and $\text{\tt\textdollar}_{k-1}$,
@@
    $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{k-i-1}$ and $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\texttt{c}$ for $i\in[0\mathinner{.\,.}k-1]$,
@@
    $\texttt{a}^{k-i-1}\texttt{c}\text{\tt\textdollar}_{k-i-2}$ for $i\in[1\mathinner{.\,.}k-2]$, and $\texttt{a}^{k-2}\texttt{c}\text{\tt\textdollar}_{k-2}$,
@@
    $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}$ for $i\in[1\mathinner{.\,.}k-1]$.
This sums to a total of $4k$ super-maximal extensions in $w_{k}^{R}$. Thus, $\mathtt{sre}(w_{k})-\mathtt{sre}(w_{k}^{R})=(5k-1)-4k=k-1$. This is $\Theta(\sqrt{|w_{k}|})$ since $|w_{k}|=2k^{2}+4k$. ∎
As the construction is quite complex, we give an example of the words $w_{k}$ and $w_{k}^{R}$ of Lemma 7, and their super-maximal right-extensions.
Example 1
Let $w_{3}=\texttt{c}\texttt{baa}\text{\tt\#}_{0}\texttt{baa}\text{\tt\textdollar}_{0}\texttt{c}\texttt{aba}\text{\tt\#}_{1}\texttt{aba}\text{\tt\textdollar}_{1}\texttt{c}\texttt{aab}\text{\tt\#}_{2}\texttt{aab}\text{\tt\textdollar}_{2}$. It can be verified that the super-maximal right-extensions of $w_{3}$ are:
@@
    $\texttt{baa}\text{\tt\#}_{0}$ and $\texttt{baa}\text{\tt\textdollar}_{0}$; $\texttt{aba}\text{\tt\#}_{1}$ and $\texttt{aba}\text{\tt\textdollar}_{1}$; $\texttt{aab}\text{\tt\#}_{2}$ and $\texttt{aab}\text{\tt\textdollar}_{2}$;
@@
    ca and cb; caa and cab;

### v5

@@ 4.1 Appending and prepending symbols
Cases 1 and 2 induce at most one new supermaximal right-extension in total for all possible $xc$, namely the longest right-extension in $wc$ that is a suffix of $wc$. For Case 3, consider a fixed $a\in\Sigma$. For all the increasing-length suffixes $x_{1}c,x_{2}c,\dots,x_{t}c$ of $wc$ that became right-extensions together with $x_{1}a,x_{2}a,\dots,x_{t}a$, one can see that the latter form a chain of suffixes of $x_{t}a$. Hence, we only have one possible new supermaximal right-extension ending with $a$, namely $x_{t}a$. Observe that the chain of suffixes $x_{1}a,x_{2}a,\dots,x_{t}a$ is unique: if the suffix $x$ is always followed by $a$, any suffix $y$ of $x$ is either right-maximal in $w$ (and $y$ falls within Case 2), or it is always followed by an $a$ (because $x$ is always followed by an $a$), i.e. $y=x_{i}$ for some $i\in[1\mathinner{.\,.}t]$. ∎
Lemma 3.
@@
Proof.
The lower bound follows from Lemma 1. For the upper bound, let cxacxa be the shortest prefix of $cw$ that is not a right-extension of $w$, but is a right-extension of $cw$ (if it exists). This implies that there exists $b\neq a$ such that $cxb\in\mathcal{F}_{w}$ and $cxa\not\in\mathcal{F}_{w}$ (otherwise, cxacxa would be a right-extension of $w)$, so no prefix of $cw$ of length $|cxa|$ or more is right-maximal, and thus no prefix longer than cxacxa can be a right-extension. By the minimality of cxacxa, all prefixes of $cw$ shorter than cxacxa either are not right-extensions of $cw$, or are right-extensions of both $w$ and $cw$. Therefore, cxacxa together with some cxbcxb appearing in $w$, are the only possible new right-extensions in $cw$ with respect to $w$. ∎
Lemmas 2 and 3 yield remarkably simple algorithms to compute smallest suffixient sets, which we detail in Section 7.
By letting $c=\text{\tt\textdollar}\not\in\mathcal{F}_{w}$ in Lemma 2, we relate $\chi$ to $\mathtt{sre}$ (note that $\chi$ is always at least $\mathtt{sre}+1$ because of the new supermaximal extension ending with \$). This makes clear the relation between Combinatorics on words [21] with suffixient sets, via the common notion of right-special factors (which we call here right-maximal substrings).
Corollary 1.
@@
4.2 Sensitivity to edit operations and rotations
We first prove some asymptotic relations between the additive sensitivities of edit operations and rotations.
Lemma 5.
$AS_{\chi,\rho}(n)=\Theta(AS_{\chi,\mathcal{R}}(n))$ for $\rho\in\{\mathtt{del},\mathtt{sub}\}$, and $AS_{\chi,\mathtt{ins}}(n)=O(AS_{\chi,\mathcal{R}}(n))$.
Proof.
We start proving $AS_{\chi,\rho}(n)=O(AS_{\chi,\mathcal{R}}(n))$, with $\rho=\mathtt{ins}$. Let $|xy|=n$ and let us insert $c\in\Sigma$ between $x$ and $y$. Let us also shorten $R(n)=AS_{\chi,\mathcal{R}}(n)$. Then it holds that
$\mathtt{sre}(xcy)$ $\leq$ $\mathtt{sre}(yxc)+R(n+1)\leq\mathtt{sre}(yx)+2+R(n+1)$
$\leq$ $\mathtt{sre}(xy)+R(n)+2+R(n+1),$
where the second inequality stems from Lemma 2. Since $1\leq R(n)=O(n)$ because $\mathtt{sre}=O(n)$, it follows that $\mathtt{sre}(xcy)=\mathtt{sre}(xy)+O(AS_{\chi,\mathcal{R}}(n))$.
For $\rho=\mathtt{del}$, we instead delete $d$ from xdyxdy, so we have
$\mathtt{sre}(xy)\leq\mathtt{sre}(yx)+R(n)\leq\mathtt{sre}(yxd)+R(n)\leq\mathtt{sre}(xdy)+R(n+1)+R(n),$
where again the second inequality stems from Lemma 2. By combining both equations above, we ge a bound for $\rho=\mathtt{sub}$, with $n=|xcy|$:
$\mathtt{sre}(xcy)\leq\mathtt{sre}(yx)+2+R(n)\leq\mathtt{sre}(yxd)+4+R(n)\leq\mathtt{sre}(xdy)+4+2R(n).$
In the other direction, for $AS_{\chi,\mathcal{R}}(n)=O(AS_{\chi,\mathrm{del}}(n))$, let \$ be a symbol not in $xy$ and $D(n)=AS_{\chi,\mathtt{del}}(n)$. Then
$\mathtt{sre}(yx)\leq\mathtt{sre}(y\text{\tt\textdollar}x)+D(n+1)\leq\mathtt{sre}(xy\text{\tt\textdollar})+D(n+1)\leq\mathtt{sre}(xy)+2+D(n+1).$
[…]
$\mathtt{sre}(ycx)$ $\leq$ $\mathtt{sre}(y\text{\tt\textdollar}x)+S(n)\leq\mathtt{sre}(xy\text{\tt\textdollar})+S(n)$
$\leq$ $\mathtt{sre}(xy)+2+S(n)\leq\mathtt{sre}(xyc)+2+S(n).\qquad\qquad\qquad\qed$
We now show that $\mathtt{sre}$ can grow by $\Omega(\sqrt{n})$ upon arbitrary edits and rotations.
Lemma 6.
Let $w_{m}=\texttt{a}\texttt{b}^{2m}\texttt{a}\texttt{b}^{2m+2}\prod_{k=1}^{m-1}u_{k}$, where $u_{k}=(\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k})^{2}$ for some $m\geq 3$. Then, $\mathtt{sre}(w_{m})=6m+4$ holds.
Proof.
Let us consider the prefix $v=\texttt{a}\texttt{b}^{2m}\texttt{a}\texttt{b}^{2m+2}$ of $w$. The substrings $\texttt{a}\texttt{b}^{2m}$ and $\texttt{b}^{2m+1}$ appear twice in $v$, once followed by a and once followed by b. For each $k\in[1,m-1]$: the substring $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}$ appears three times, once in $v$, and twice in $u_{k}$ (note that one occurrence of $\texttt{b}^{m-1}\texttt{a}\texttt{b}^{m+1}$ is a suffix of $w_{m}$); the substring $\texttt{b}^{2m-k+1}\texttt{a}\texttt{b}^{k}$ appears once in $v$ and twice across $u_{k}u_{k+1}$; the substring $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}$ appears twice across $u_{k-1}u_{k}u_{k+1}$.
By carefully analyzing the right-extensions of these factors, one can verify that the following substrings are supermaximal right-extensions of $w_{m}$:
@@ 4.2 Sensitivity to edit operations and rotations
    $\texttt{a}\texttt{b}^{2m}\texttt{a}$ and $\texttt{a}\texttt{b}^{2m+1}$,
@@
    $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}$ and $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k+1}$ (for $k\in[1,m-1]$),
@@
    $\texttt{b}^{2m-k+1}\texttt{a}\texttt{b}^{k}\texttt{a}$ and $\texttt{b}^{2m-k+1}\texttt{a}\texttt{b}^{k+1}$ (for $k\in[1,m-1]$),
@@
    $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}$ and $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k+1}$ (for $k\in[1,m-2]$),
  5. 5.
    $\texttt{b}^{2m+1}\texttt{a}$ and $\texttt{b}^{2m+2}$.
We now prove that there are no other super-maximal right-extensions in $w$. Let us consider the other right-extensions (also see Figure 2). For any other right-extensions $x$ ending in a, we have that it is a suffix of one of the supermaximal right-extensions, since every occurrence of a, except the first and last a, serves as the ending position of a supermaximal right-extension listed above; the longest right-extensions ending in correspondence of the first and last a are a and $\texttt{b}^{m+1}\texttt{a}\texttt{b}^{m-1}\texttt{a}$ respectively, and both are suffixes of $\texttt{b}^{m+2}\texttt{a}\texttt{b}^{m-1}\texttt{a}$ (case 3. with k=m1k=m-1). For any other right-extension ending in b: by construction, any substring including a at least twice is not a right-extension, with the exception of $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k+1}$ for $k\in[1,m-2]$ (included in case 4.); all right-extensions containing only one a have the form $\texttt{b}^{i}\texttt{a}\texttt{b}^{j}$, for any $i\geq 0$$j\geq 1$ such that $i+j\leq 2m+1$ and $j\neq m$, and each is a suffix of one of the supermaximal right-extensions included in either cases 2. or 3.; finally, the right-extensions that do not contain a’s consist of runs of b’s, which are all suffixes of $\texttt{b}^{2m+2}$ (included in 5.). Therefore, there are no other supermaximal extensions, and $\mathtt{sre}(w_{m})=6m-4$. ∎
Figure 2: The supermaximal right-extension of the string $w_{4}$ shown in Example 1. Note that the ending positions of those ending within $\texttt{b}^{10}$ are all different.
Example 1.
Let $w_{4}=\texttt{a}\texttt{b}^{8}\texttt{a}\texttt{b}^{10}\texttt{a}\texttt{b}\texttt{a}\texttt{b}^{7}\texttt{a}\texttt{b}\texttt{a}\texttt{b}^{7}\texttt{a}\texttt{b}^{2}\texttt{a}\texttt{b}^{6}\texttt{a}\texttt{b}^{2}\texttt{a}\texttt{b}^{6}\texttt{a}\texttt{b}^{3}\texttt{a}\texttt{b}^{5}\texttt{a}\texttt{b}^{3}\texttt{a}\texttt{b}^{5}$ (also see Figure 2). It can be verified that the supermaximal right-extensions of $w_{4}$ are:
@@
    $\texttt{a}\texttt{b}^{8}\texttt{a}$ and $\texttt{a}\texttt{b}^{9}$,
@@
    $\texttt{b}\texttt{a}\texttt{b}^{7}\texttt{a}$ and $\texttt{b}\texttt{a}\texttt{b}^{8}$; $\texttt{b}^{2}\texttt{a}\texttt{b}^{6}\texttt{a}$ and $\texttt{b}^{2}\texttt{a}\texttt{b}^{7}$; $\texttt{b}^{3}\texttt{a}\texttt{b}^{5}\texttt{a}$ and $\texttt{b}^{3}\texttt{a}\texttt{b}^{6}$,
@@
    $\texttt{b}^{8}\texttt{a}\texttt{b}\texttt{a}$ and $\texttt{b}^{8}\texttt{a}\texttt{b}^{2}$; $\texttt{b}^{7}\texttt{a}\texttt{b}^{2}\texttt{a}$ and $\texttt{b}^{7}\texttt{a}\texttt{b}^{3}$; $\texttt{b}^{6}\texttt{a}\texttt{b}^{3}\texttt{a}$ and $\texttt{b}^{6}\texttt{a}\texttt{b}^{4}$,
@@
    $\texttt{b}^{7}\texttt{a}\texttt{b}\texttt{a}\texttt{b}^{7}\texttt{a}\texttt{b}\texttt{a}$ and $\texttt{b}^{7}\texttt{a}\texttt{b}\texttt{a}\texttt{b}^{7}\texttt{a}\texttt{b}^{2}$; $\texttt{b}^{6}\texttt{a}\texttt{b}^{2}\texttt{a}\texttt{b}^{6}\texttt{a}\texttt{b}^{2}\texttt{a}$ and $\texttt{b}^{6}\texttt{a}\texttt{b}^{2}\texttt{a}\texttt{b}^{6}\texttt{a}\texttt{b}^{3}$,
  5. 5.
    $\texttt{b}^{9}\texttt{a}$ and $\texttt{b}^{10}$.
One can see that $\mathtt{sre}(w_{4})=20$, as stated in Lemma 6.
Lemma 7.
Let $w_{m}=\texttt{a}\texttt{b}^{2m}\texttt{a}\texttt{b}^{2m+2}\prod_{k=1}^{m-1}u_{k}$, where $u_{k}=(\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k})^{2}$ for some $m\geq 3$. It holds:
  1. (Ins)
    If $w_{m}^{\prime}=\texttt{a}\texttt{b}^{4m+2}\prod_{k=1}^{m-1}u_{k}$, then $\mathtt{sre}(w_{m}^{\prime})=4m-2$;
  2. (Sub)
    If $w_{m}^{\prime}=\texttt{a}\texttt{b}^{4m+3}\prod_{k=1}^{m-1}u_{k}$, then $\mathtt{sre}(w_{m}^{\prime})=4m-2$;
  3. (Del)
    If $w_{m}^{\prime}=\texttt{a}\texttt{b}^{2m}\texttt{a}\texttt{a}\texttt{b}^{2m+2}\prod_{k=1}^{m-1}u_{k}$, then $\mathtt{sre}(w_{m}^{\prime})=4m+2$;
  4. (Rot)
    If $w_{m}^{\prime}=\texttt{a}\texttt{b}^{2m+2}\prod_{k=1}^{m-1}u_{k}\cdot\texttt{a}\texttt{b}^{2m}$, then $\mathtt{sre}(w_{m}^{\prime})=4m-2$.
Proof.
We prove the claims separately.
Insertions. Let us consider the string $w_{m}^{\prime}$ obtained by removing the second a from $w_{m}$. Then, the prefix $v=\texttt{a}\texttt{b}^{2m}\texttt{a}\texttt{b}^{2m+2}$ of $w_{m}$ is changed to $v^{\prime}=\texttt{a}\texttt{b}^{4m+2}$. The substring $\texttt{b}^{2m-k+1}\texttt{a}\texttt{b}^{k}$ occurs only once for each $k\in[1..m-1]$, and any subtring $\texttt{b}^{i}\texttt{a}\texttt{b}^{j}$ of $w_{m}$ with $i+j>2m$ does not occur in $w^{\prime}_{m}$. Instead, $\texttt{b}^{4m+1}$ occurs twice in $v^{\prime}$, and $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k-1}\texttt{a}$ and $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}$ become new supermaximal right-extensions for each $k\in[1,m-2]$, occurring twice in $u_{k+1}$ and $u_{k}$, respectively. Also, $\texttt{a}\texttt{b}^{2m-1}$ appears once in $v^{\prime}$ and twice in $u_{1}$, and $\texttt{b}^{m-1}\texttt{a}\texttt{b}^{m-1}$ can be found three times in $u_{m-1}$. Wrapping up and analyzing its right-extensions, one can verify that the supermaximal right-extensions of $w_{m}^{\prime}$ and their ending positions are the following:
@@
    $\texttt{a}\texttt{b}^{2m-1}\texttt{a}$ (ending at $(2m+3)$ in $v^{\prime}$), and $\texttt{a}\texttt{b}^{2m}$ (ending at $(2m+3)$ in $u_{1}$ and $1$ in $u_{2}$),
@@
    $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k-1}\texttt{a}$ (ending at $(2m+3)$ in $u_{k+1}$ and $1$ in $u_{k+2}$) and $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}$ (ending at $(2m+2)$ and $(4m+4)$ in $u_{k}$) for $k\in[1,m-2]$,
@@
    $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}$ (ending at $(2m+k+4)$ in $u_{k}$)and $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k+1}$ (ending at $(k+2)$ in $u_{k+1}$) for $k\in[1,m-2]$,
@@
    $\texttt{b}^{4m+1}\texttt{a}$ (ending at $(4m+4)$ in $v^{\prime}$) and $\texttt{b}^{4m+2}$ (ending at $(4m+3)$ in $v^{\prime}$),
  5. 5.
    $\texttt{b}^{m-1}\texttt{a}\texttt{b}^{m-1}\texttt{a}$ (ending at $(2m+1)$ in $u_{m-1}$) and $\texttt{b}^{m-1}\texttt{a}\texttt{b}^{m}$ (ending at $(m+1)$ and $(3m+3)$ in $u_{m-1}$) .
Thus, we have $\mathtt{sre}(w_{m}^{\prime})=4m-2$.
Substitutions. Assuming that the string $w_{m}^{\prime}$ is obtained by replacing the second a in $w_{m}$ with b, we obtain a string similar to the one we consider in case Insertions. Considering its structure, we have that the supermaximal right-extensions of $w_{m}^{\prime}$ are the following:
@@
    $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k-1}\texttt{a}$ and $\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}$ (for $k\in[1,m-2]$),
@@
    $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}$ and $\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k}\texttt{a}\texttt{b}^{2m-k}\texttt{a}\texttt{b}^{k+1}$ (for $k\in[1,m-2]$),